OpenStudy (anonymous):

I GIVE MEDALS!!!!!!!! Solve the inequality. Explain all steps. 3a – 4 ≤ 5

OpenStudy (akashdeepdeb):

3a - 4 <= 5 3a - 4 + 4 <= 5+4 3a <= 9 a <= 9/3 a <= 3 Understood? :)

OpenStudy (anonymous):

Somewhat!

OpenStudy (akashdeepdeb):

3a - 4 <= 5 3a - 4 + 4 <= 5+4 [Add +4 both sides to cancel off 4 with -4 in the LHS] 3a <= 9 [Divide both sides by 3 to get the value of 3] a <= 9/3 a <= 3 Understood now? :)

OpenStudy (anonymous):

Oh okay thanks!

OpenStudy (akashdeepdeb):

:)